# Inheritune: Training Smaller Yet More Attentive Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, primarily due to the transformer architecture and its self-attention mechanism. However, we observe that in standard decoder-style LLMs, attention matrices degenerate to single-column for deeper layers. Layers in this state are unable to learn anything meaningful and mostly redundant; we refer to these as lazy layers. The goal of this paper is to train smaller models by eliminating this structural inefficiency without compromising performance.
Motivated by this observation, we propose Inheritune, a simple yet effective training recipe for developing smaller, high-performing language models. Smaller models trained with $\mathsf{Inheritune}$ inherit early transformer layers from a larger pre-trained model, then retrain and progressively expand until they match or exceed the performance of the larger model. We demonstrate that $\mathsf{Inheritune}$ enables the training of various sizes of GPT-2 models on datasets like OpenWebText-9B and FineWeb_Edu. Models trained with $\mathsf{Inheritune}$, despite having significantly fewer layers, match or even surpass the performance of their larger counterparts. For instance, our 16-layer GPT-2 medium variant achieves comparable performance to the standard 24-layer GPT-2 medium model. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
@@ 1 Introduction
Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017). These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked together, building model capacity. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data. As these models scale in depth and size, their performance tends to improve Kaplan et al. (2020); Hoffmann et al. (2022), benefiting from increased model capacity.
@@
The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block. It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data. However, as models grow in depth, they often encounter a phenomenon known as attention degeneration, caused by a collapse in the attention rank Noci et al. (2022); Dong et al. (2021); He et al. (2023). Notably, this phenomenon has not been studied in the context of standard LLMs. A formal discussion on attention degeneration is provided in Section 2.
@@
In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al. (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices. Further investigation reveals that most of these rank-1 matrices are also single-column, i.e., their mass is concentrated in a single column. Our attention matrix analysis is presented in Figure 1. We term these deeper layers, where all attention matrices of a given layer are degenerate, as lazy layers.
@@
Motivated by this new finding, we aim to develop performant small base language models (LMs) that utilize weights from inefficient larger base LMs. A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple, yet effective, training approach as $\mathsf{Inheritune}$.
(a) Rank analysis of vanilla GPT-2 Medium
(b) Matrix mass analysis of vanilla GPT-2 Medium
(c) Model initialization with later layers
(d) Rank analysis of vanilla GPT-2 Large
(e) Matrix mass analysis of vanilla GPT-2 Large
(f) Model initialization with later layers
Figure 1: Attention matrices of many deeper layers often degenerate to single-column matrices in regular decoder style LLMs, layers with fully degenerate attention fail to learn meaningful representations. We computed a single attention matrix with 100 tokens from the OpenWebText validation set with 4M tokens. Next we performed 100 runs and plotted the mean and std of the max rank and mass as a function of layers for our rank and mass analysis respectively. Figure (a and d): An analysis of a 24-layer GPT-2 medium and a 36-layer GPT-2 large shows the max rank of the attention matrices across all layers. Figure (b and e): A closer look at the same GPT-2 models also reveals that the dominant mass proportion of several attention matrices is concentrated in a single-column, particularly in deeper layers. Figure (c and f): When initializing 12-layer and 18-layer variants of the vanilla GPT-2 medium and GPT-2 large models with deeper layers (Lazy layers) exhibiting degenerated attention–their performance is comparable to models with random initialization. However, initializing models with early layers leads to significantly better generalization and convergence.
@@
    Analysis of Attention Degeneration Leading to Lazy Layers. We empirically investigate attention degeneration in standard LLM settings. Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (see Figure 1 and Figure 2).
@@
    Introduction of Inheritune. We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models. This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model. The initialization is entirely zero-shot. We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText dataset with 9B tokens (with data repetition) and the FineWeb_edu dataset with 100B tokens (without data repetition).
@@
    Evaluation Against Multiple Baselines. Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Table 1), model initialization and efficient training baselines (refer Table 2), and models trained using two knowledge distillation techniques (refer Figure 3). In settings where training tokens are not repeated, we observe similar trends (refer Figure 4).
L4 H2
L7 H3
L20 H2
L22 H4
24 layer GPT2 medium model trained from scratch.
L4 H2
L7 H3
L11 H2
L15 H1
Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
Figure 2: Inheritune preserves effective attention patterns in smaller models. Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected strings, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s later layers.
2 Attention Degeneration in Standard Decoder-Style LLMs
@@
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times d}$, where $T$ denotes the sequence length (number of tokens), and $d$ represents the embedding dimension or hidden size of the model. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times d}$.
Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN). The self-attention mechanism enables the model to weigh the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as
$\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$ (1)
@@
The attention matrix $A(X)\in\mathbb{R}^{T\times T}$ captures the pairwise attention scores between all token positions in the sequence. The softmax is applied row-wise. The attention matrix $A(X)$ is then used to compute a weighted sum of the value vectors.
Previous research by Dong et al. (2021) and He et al. (2023) has shown that in self-attention networks (SANs) without residual connections and feed-forward networks (FFNs), the rank of an attention matrix converges to rank-1 doubly exponentially with respect to the depth of the model. This phenomenon, known as the rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly. Noci et al. (2022) showed that even with residual connections (without layernorm) attention matrices can still lose rank in deeper layers if the residual connections are not scaled by $1/\sqrt{L}$. Interestingly, they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers, which affects the overall trainability of the transformer-based models. However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layer norms, and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
@@
In this paper, we deeply analyzed the structure of attention matrices to diagnose the presence of rank collapse or similar phenomena in standard transformer-based LLMs using GPT-2 models. For our first analysis, we compute the approximate rank (referred to as rank hereafter) of $A(X)$ for all attention heads within each layer. Formally, we began by computing the Singular Value Decomposition (SVD) of $A(X)=U\Sigma V^{\top}$, where the diagonal entries $\{\sigma_{i}\}_{i=1}^{T}$ of $\Sigma$ represent the singular values, quantifying the variance captured by each corresponding singular vector of $A(X)$. To determine the minimal number of singular values required to capture 90% of the total variance, we solved: $k^{*}=\min\left\{k\in\{1,2,\ldots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq 0.90\right\}$. Here, $k^{*}$ is the approximate rank of $A(X)$ computed using the explained variance method.
Dominant Single-Column Structure in Attention Matrices
We further investigated the dominant structure of these rank-1 attention matrices and observed that, on average, many of these matrices have their mass concentrated in a single column. This intrinsic structure can be viewed as a special case of rank-1 attention matrices. To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$. Next, to determine the minimal number of columns required to capture 90% of the total mass, we solved; $m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq 0.90\right\}$. Here, $m^{*}$ represents the minimal number of columns needed for the cumulative column mass ratios to reach or exceed 90%.
Degeneration of Attention Matrices in GPT-2 Models
@@
In Figure 1, we present the layer-wise analysis of rank and mass. For this analysis, we computed $A(X)$ using 100 randomly selected samples from the validation set of OpenWebText valset with 4M tokens, each with a sequence length of $T=100$ tokens, across all attention heads within each layer. Next, we computed the rank with a 90% variance threshold, and for every layer, we chose the maximum rank across all the heads. In Figure 1(a) and 1(d) we plotted maximum rank as a function of layers for 100 runs with mean and standard deviation (std); it is quite evident that many deeper layers exhibit all rank-1 attention matrices. We highlight that this rank collapse occurs for both the GPT-2 medium and GPT-2 Large models with skip connections and FFNs, extending the findings of Dong et al. (2021) and Noci et al. (2022) to standard LLM architectures. Next, based on our mass analysis, we demonstrate that most of the rank collapsed attention matrices are also single-column matrices as depicted in Figure 1(b) and Figure 1(e). We follow the analysis protocol as discussed in the rank analysis. This degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit complete degeneration of all attention matrices across all attention heads, indicating reduced performance and less effective token mixing.
@@
How much transferable knowledge do lazy layers hold compared to their early counterparts? To answer this question, we initialized a 12-layer GPT-2 medium variant [^1] A variant shares the same configurations as the parent model but has fewer layers. and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models. These pre-trained models are trained on the OpenWebText-9B dataset for 100K steps. We then fine-tuned these GPT-2 variants on the same dataset for an additional 10K steps. For comparison, we conducted two baseline experiments in which the GPT-2 variants were initialized either with the first few transformer blocks or with random initialization. As shown in Figures 1(c) and 1(f), models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provides additional evidence that lazy layers with fully degenerate attention fail to learn meaningful representations.
2.1 Attention Pattern Visualization
To provide further evidence of lazy layers and provide a preview of our solution, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model. Fig. 2 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. We visualized full attention and not causal attention for the sake of better visualization. In practice, GPT-2 models compute causal attention. We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over three runs.
Algorithm 1 Inheritune: Training Recipe for Small Language Models
1:Reference model $\mathcal{M}_{\text{ref}}$ with $k$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
2:Initialize $\mathcal{M}_{\text{tgt}}$ with first $n=k/2$ layers from $\mathcal{M}_{\text{ref}}$
3:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
4:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
5:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
6:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
7:end while
8:return Optimized model $\mathcal{M}_{\text{tgt}}$
In the 24-layer model trained from scratch (top row of Fig. 2), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al. (2019); Beltagy et al. (2020); Chen et al. (2021). In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus. This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes the model more attentive and addresses attention degeneration, potentially leading to more efficient models in a compact size (also refer to Figure 9). We will discuss these results in more detail after introducing our method, but this preview underscores the promise of our approach.

### v3

# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

@@ Abstract
Large Language Models (LLMs) leverage the transformer architecture and its self-attention mechanism to achieve remarkable performance. However, we identify a structural inefficiency in standard pre-trained decoder-style LLMs: attention matrices in many deeper layers frequently degenerate, often collapsing to near rank-one, single-column structures. These layers, which we term lazy layers, are redundant and inefficient.
To address this, we introduce $\mathsf{Inheritune}$, a simple yet effective training recipe for developing smaller, more attentive, and high-performing language models. $\mathsf{Inheritune}$ initializes a compact model by inheriting the potent early layers from a larger pre-trained model, then progressively retrains and expands it. Our demonstration of $\mathsf{Inheritune}$’s efficacy on various models and datasets reveals that models trained with this approach, despite having significantly fewer layers, can achieve performance comparable to or even exceeding their larger counterparts. This approach enables model compression by design, producing language models that are both compact and high-performing. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
@@ 1 Introduction
Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017). These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked with eachother building model capacity. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data. As these models scale in depth and size, their performance tends to improve Kaplan et al. (2020); Hoffmann et al. (2022), benefiting from increased model capacity.
@@
The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block. It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data. However, as models grow in depth, they often encounter a phenomenon known as attention degeneration caused by collapse in the attention rank ((Noci et al., 2022; Dong et al., 2021; He et al., 2023)). Notably, this phenomenon has not been studied in the context of standard LLMs. A formal discussion on attention degeneration is provided in Section 2.
@@
In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al. (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices. Further investigation reveals that most of these rank-1 matrices are also single-column, i.e. their mass is concentrated to a single column. Our attention matrix analysis is presented in Figure 1. We term these deeper layers, where all attention matrices of a given layer are degenerated, as lazy layers.
@@
Motivated by this new finding we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs. A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple yet effective training approach as $\mathsf{Inheritune}$.
(a) GPT-2 Medium (355M) L=24, H=1024
(b) GPT-2 Large (770M) L=36, H=1280
(c) GPT-2 XLarge (1.5B) L=48, H=1600
Figure 1: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near rank-1, limiting their ability to learn meaningful representations. We compute $\text{MaxRank}^{(l)}$ (averaged over $N=100$ randomly selected sequences each with $T=100$ tokens) for each layer $l$ using the OpenWebText validation set. Our rank analysis of 24-layer GPT-2 medium, 36-layer GPT-2 large, and 48-layer GPT-2 xlarge models reveals that attention matrices in many deeper layers collapse to near rank-1.
@@
    Analysis of Attention Degeneration Leading to Lazy Layers. We empirically investigate attention degeneration in standard LLM settings. Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1).
@@
    Introduction of Inheritune. We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models. This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model. The initialization is entirely zero-shot. We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText and the FineWeb_edu dataset.
@@
    Evaluation Against Multiple Baselines. Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Figure 5 and Figure 6), model initialization and efficient training baselines (refer Table 1), and models trained using two knowledge distillation techniques (refer Figure 7).
2 Attention Collapse and the Emergence of Lazy Layers in LLMs
@@
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times e}$, where $T$ denotes the sequence length (number of tokens), and $e$ represents the embedding dimension or model hidden size. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times e}$.
Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN). The self-attention mechanism enables the model to weight the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as $\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$
@@
The attention matrix $A(X)\in\mathbb{R}^{T\times T}$ captures the pairwise attention scores between all token positions in the sequence. The softmax is applied row-wise. The attention matrix $A(X)$ is then used to compute a weighted sum of the value vectors. Attention rank collapse refers to the phenomenon where the attention matrices $A(X)$ of individual heads in many layers of transformer-based language models lose their expressive capacity, converging towards lower effective rank structures. Specifically, the effective rank of attention matrices significantly reduces, often approaching rank-1, limiting the model’s ability to meaningfully differentiate between token interactions across positions in the sequence. Previous research by Dong et al. (2021) and He et al. (2023) has shown that in self-attention networks (SANs) without residual connections and feed-forward networks (FFNs), the rank of an attention matrix converges to rank-1 doubly exponentially with respect to the depth of the model. This phenomenon, known as rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly. Noci et al. (2022) showed that even with residual connections (without layernorm) attention matrices can still lose rank in deeper layers if the residual connections are not scaled by $1/\sqrt{L}$. Interestingly they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers which affects the overall trainability of the transformer based models. However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layernorms and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
@@
To assess the presence and severity of rank collapse within standard decoder style transformer architectures (e.g., GPT-2, LLaMA etc.), we utilize singular value decomposition (SVD) for each attention matrix $A(X)=U\Sigma V^{\top},$ where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{T}\geq 0$. The approximate rank (referred to as rank hereafter) of an attention matrix, parameterized by a variance threshold $\tau$, is formally computed as:
$k^{*}=\min\left\{k\in\{1,2,\dots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq\tau\right\},$
where $\tau\in(0,1)$ represents the proportion of variance that must be captured by the top $k$ singular values. A lower value of $k^{*}$ indicates stronger rank collapse. In this work, we set $\tau=0.90$. A lower $k^{*}$ indicates stronger rank collapse.
@@
In Figure 1, we present the layer-wise analysis of rank of GPT-2 models. For this analysis, we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText with 4M tokens, each with a sequence length of $T=100$ tokens across all attention heads within each layer. We then define the average approximate rank for each head and layer as $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$. Subsequently, we aggregate this metric per layer by taking the maximum rank across heads: $\mathrm{MaxRank}^{(l)}=\max_{h}\{\mathrm{Rank}^{(h,l)}\}$. As demonstrated in Figure 1, $\text{MaxRank}^{(l)}$ reveals that many deeper layers exhibit attention matrices that are predominantly near rank-1. We highlight that this rank collapse occurs in GPT-2 medium, Large, and xLarge models—which are standard architectures incorporating skip connections and FFNs—thereby extending the findings of Dong et al. (2021) and Noci et al. (2022) to these widely-used LLMs. The rank analysis is quite robust to the dataset in Figure 10, we performed the same analysis using the same models on held out data from Fineweb_edu dataset (Penedo et al., 2024). Overall, the degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit a near-complete rank collapse of attention matrices across all heads, suggesting potentially reduced representational capacity and less effective token mixing in these layers.
2.1 The Functional Ineffectiveness of Lazy Layers
Having identified lazy layers, we investigate their practical utility: Do these structurally degenerated layers retain transferable knowledge, or are they functionally impaired? Our experiments suggest the latter.
(a) Rank profile (GPT-2 Small 12L).
(b) Performance of 4L variants.
Figure 2: Higher-rank layers transfer better. (Left, 2(a)) Layer-wise $\mathrm{MaxRank}^{(l)}$ of a pre-trained 12L GPT-2 Small. (Right, 2(b)) Validation loss of 4L variants initialized with potent blocks (AvgRank $\approx 8.4-9.5$) vs. a lazy block (AvgRank $\approx 1.2$) or random weights, after 100K steps. Lazy block initialization mirrors random.
In the first set of experiments, we trained a vanilla GPT2-small (125M) model with 12 layers for 100K steps on the OpenWebText dataset. We then performed the rank analysis described earlier, with results presented in Figure 2. Specifically, we aggregated the approximate ranks over groups of contiguous layers using $\text{AvgRank}=\frac{1}{L}\sum_{l=1}^{L}\text{MaxRank}^{(l)}$, where $L$ is the number of layers in each group. Subsequently, we trained three GPT2-small variants [^1] A variant shares the same configurations as the parent model but has fewer layers. for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT2-small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$. The last model is initialized with lazy layers. For comparison, we also trained another GPT2-small variant with random initialization for 100K steps. All models were trained on the OpenWebText dataset. As shown in Figure 2, the model initialized with layers from the vanilla GPT2-small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss). Additionally, we observed that the model initialized with lazy layers performed very similarly to the model with random initialization suggesting that lazy layers contain minimal transferable knowledge. The key results are also summarized in Table 4.
(a) GPT2-medium (12-layer) variants.
(b) GPT2-large (16-layer) variants.
Figure 3: When initializing 12-layer and 16-layer variants of GPT2-medium and GPT2-large with deeper (lazy) layers showing degenerated attention, performance is comparable to random initialization. In contrast, early-layer initialization leads to significantly better convergence and generalization.
@@ 2.1 The Functional Ineffectiveness of Lazy Layers
For the second set of experiments we utilized larger models namely GPT-2 medium and GPT-2 large both similarly trained for 100K steps using OpenWebText dataset. Here we initialized a 12-layer GPT-2 medium variant and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models. We then trained these GPT-2 variants on the same dataset for 10K steps. For comparison, we conducted two baseline experiments where the GPT-2 variants were initialized either with the first half of transformer layers (potent layers with high AvgRank) and with random initialization. As shown in Figure 3, models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provide additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.